ID |
Date |
Author |
Type |
Category |
Subject |
13749
|
Thu Apr 12 18:12:49 2018 |
gautam | Update | ALS | NPRO channels hijacked |
Summary:
- Today, the measured IR ALS noise for the X arm was dramatically improved. The main change was that I improved the alignment of the PSL pickoff beam into its fiber coupler.
- The noise level was non-stationary, leading me to suspect power modulation of the RF beat amplitude.
- I am now measuring the stability of the power in the two polarizations coming from EX table to the PSL table, using the PSL diagnostic connector channels.
- The EX beam is S-polarized when it is coupled into the fiber. The PSL beam is P-polarized. However, it looks like I have coupled light along orthogonal axes into the fiber, such that when the EX light gets to the PSL table, most of it is in the P-polarization, as judged by my PER measurement setup (i.e. the alignment keys at the PSL table and at the EX table are orthogonal). So it still seems like there is something to be gained by trying to improve the PER a bit more.
Details:
Today, I decided to check the power coupled into the PSL fiber for the BeatMouth. Surprisingly, it was only 200uW, while I had ~3.15mW going into it in January. Presumably some alignment drifting happened. So I re-aligned the beam into the fiber using the steering mirror immediately before the fiber coupler. I managed to get ~2.9mW in without much effort, and I figured this is sufficient for a first pass, so I didn't try too much more. I then tried making an ALS beat spectrum measurement (arm locked to IMC length using POX, green following the arm using end PDH servo). Surprisingly, the noise performannce was almost as good as the reference! See Attachment #1, in which the red curve is an IR beat (while all others are green beats). The Y arm green beat performance isn't stellar, but one problem at a time. Moreover, the kind of coherence structure between the arm error signal and the ALS beat signal that I reported here was totally absent today.
Upon further investigation, I found that the noise level was actually breathing quite significantly on timescales of minutes. While I was able to successfully keep the TEM00 mode of the PSL beam resonant inside the arm cavity by using the ALS beat frequency as an error signal and MC2 as a frequency actuator, the DC stability was very poor and TRX was wandering around by 50%. So my new hypothesis is that the excess ALS noise is because of one or more of
- Beam jitter at coupling point into fiber.
- Polarization drift of the IR beams.
While I did some work in trying to align the PSL IR pickoff into the fiber along the fast (P-pol) axis, I haven't done anything for the X end pickoff beam. So perhaps the fluctuations in the EX IR power is causing beatnote amplitude fluctuations. In the delay line + phase tracker frequency discriminator, I think RF beatnote amplitude fluctuations can couple into phase noise linearly. For such an apparently important noise source, I can't remeber ever including it in any of the ALS noise budgets.
Before Ph237 today, I decided to use my polarization monitoring setup and check the "RIN" of power in the two polarizations coming out of the fiber on the PSL table. For this purpose, I decided to hijack the Acromag channels used for the PSL diagnostics connector Attachment #2 shows that there is fluctuations at the level of ~10% in the p-polarization. This is the "desired" polarization in that I aligned the PSL beam into the fiber to maximize the power in this polarization. So assuming the power fluctuations in the PSL beam are negligible, this translates to sqrt(10) ~3% fluctuation in the RF beat amplitude. This is at best a conservative estimate, as in reality, there is probably more AM because of the non PM fibers inside the beatmouth.
All of this still doesn't explain the coherence between the measured ALS noise and the arm error signal at 100s of Hz (which presumably can only happen via frequency noise in the PSL).
Another "mystery" - yesterday, while I was working on recovering the Y arm green beat signal on the PSL table, I eventually got a beat signal that was ~20mVpp into 50ohms, which is approximately the same as what I measured when the Y arm ALS performance was "nominal", more than a year ago. But while viewing the Y arm beats (green and IR) simultaneously on an o'scope, I wasn't able to keep both signals synchronised while triggering on one (even though the IR beat frequency was half the green beat frequency). This means there is a huge amount of relative phase noise between the green and IR beats. What (if anything) does this mean? The differential noise between these two beats should be (i) phase noise at the fiber coupler/ inside the fiber itself, and (ii) scatter noise in the green light transmitted through the cavity. Is it "expected" that the relative phase noise between these two signals is so large that we can't view both of them on a common trigger signal on an o'scope? Also - the green mode-matching into the Y arm is abysmal.
Anyways - I'm going to try and tweak the PER and mode-matching into the X end fiber a little and monitor the polarization stability (nothing too invasive for now, eventually, I want to install the new fiber couplers I acquired but for now I'll only change alignment into and rotation of the fiber coupler on the EX table). It would also be interesting to compare my "optimized" PSL drift to the unoptimized EX power drift. So the PSL diagnostic channels will not show any actual PSL diagnostic information until I plug it back in. But I suspect that the EPICS record names and physical channel wiring are wrong anyways - I hooked up my two photodiode signals into what I would believe is the "Diode 1 Power" and "Laser crystal temperature" monitors (as per the schematic), but the signals actually show up for me in "Diode 2 Power" (p-pol) and "Didoe 1 Temperature" (s-pol).
Annoyingly, there is no wiring diagram - on my todo list i guess...
@Steve - could you please take a photo of the EX table and update the wiki? I think the photo we have is a bit dated, the fiber coupler and transmon PDs aren't in it... |
Attachment 1: IR_ALS_20180412.pdf
|
|
Attachment 2: BeatMouthDrift.png
|
|
Attachment 3: ETMX_20180416.jpg
|
|
13748
|
Thu Apr 12 10:15:33 2018 |
Kira | Update | PEM | MEDM setup |
Another update. I've changed the on/off button so that it's visible which state it's in. I did that by changing the type of C1:PEM-SEIS-EX_TEMP_SLOWLOOP from ai to bi (I checked the FSS script and copied the entry for the slowloop). Previously, MEDM was giving me an error that it wasn't an ENUM value when I wanted to use a choice button to indicate the value of slowloop, and this solved the issue. I've also added a StripTool button. |
Attachment 1: MEDM_3.png
|
|
13747
|
Wed Apr 11 10:47:26 2018 |
Steve | Update | SUS | satellite amps labeled |
Satellite amplifiers labeled with date. Old labels left on. |
Attachment 1: DSC00912.JPG
|
|
13746
|
Wed Apr 11 01:34:31 2018 |
gautam | Update | IOO | Activities today |
[kevin, gautam]
activities done today - details/plots/data/evidence tomorrow.
- Checked XARM loop shape. Updated NB code to fetch POX data from NDS and undo loop shape rather than using calibration filter bank.
- Checked POX loop calibration (m/ct). Number I was using was 8e-13. Tonight we measured 9e-13. Updated filter bank.
- Tried to get Y arm green ALS going.
- Improved GTRY from ~0.05 to 0.3.
- Tried to improve mode matching onto BBPD on PSL table to see a green beat.
- But we were unsuccessful.
- I think I got the near and far field alignment right, and the EY laser temp is set such that I can see an IR beat @~30MHz (so green beat should be at 60 MHz).
- But I couldn't see anything with scope or with HP spec analyzer.
- More tomorrow. Motivation to get green ALS working is to get some more confidence that the excess noise is indeed on the PSL light.
|
13745
|
Tue Apr 10 15:42:08 2018 |
Kira | Update | PEM | MEDM setup |
An update to the screen. I changed the min/max values for some of the parameters, as well as changing the script so that I could specify the integral gain in terms of 1e-5. I've also added this screen to the PEM tab in the sitemap. |
Attachment 1: MEDM_2.png
|
|
13744
|
Tue Apr 10 14:28:44 2018 |
gautam | Update | IOO | Further debugging |
I am working on IMC electronics. IMC is misaligned until further notice. |
13743
|
Mon Apr 9 23:49:50 2018 |
rana | Update | PSL | PSL chans |
I think we can scrap the 126MOPA channels since they're associated with the Lightwave NPRO and MOPA. We should add the channels that we need for monitoring the Innolight NPRO from the d-sub connector on its controller. |
13742
|
Mon Apr 9 23:28:49 2018 |
johannes | Configuration | DAQ | c1psl channel list |
I made a list of all the physical c1psl channels to get a better idea for how many acromags we need to replace it eventually. There 3123 unit is the one whose failure had prevented c1psl from booting, which is why it was unplugged (elog post 12852), and its channels have been inactive since. Are the 126MOPA channels used for the current mephisto? 126 tells me it's for an old lightwave laser, but I was checking a few and found that they have non-zero, changing values, so they may have been rewired.
It also hosts some virtual channels for the ISS with root C1:PSL-ISS_ defined in iss.db and dc.db, the PSL particle counter with root C1:PEM- defined in PCount.db and a whole lot of PSL status channels defined in pslstatus.db. Transfering these virtual channels to a different machine is almost trivial, but the serial readout of the particle counter would have to find a new home.
Long story short - we need:
Function |
Type |
# Channels |
#Channels (no MOPA) |
# Units |
# Units (no MOPA) |
ADC |
XT1221 |
34 |
21 |
5 |
3 |
DAC |
XT1541 |
17 |
14 |
3 |
2 |
BIO |
XT1111 |
19 |
10 |
2 |
1 |
3113 - ADC
C1:PSL-126MOPA_126PWR
C1:PSL-126MOPA_DTMP
C1:PSL-126MOPA_LTMP
C1:PSL-126MOPA_DMON
C1:PSL-126MOPA_LMON
C1:PSL-126MOPA_CURMON
C1:PSL-126MOPA_DTEC
C1:PSL-126MOPA_LTEC
C1:PSL-126MOPA_CURMON2
C1:PSL-126MOPA_HTEMP
C1:PSL-126MOPA_HTEMPSET
C1:PSL-FSS_RFPDDC
C1:PSL-FSS_LODET
C1:PSL-FSS_FAST
C1:PSL-FSS_PCDRIVE
C1:PSL-FSS_MODET
C1:PSL-FSS_VCODETPWR
C1:PSL-FSS_TIDALOUT
C1:PSL-PMC_RFPDDC
C1:PSL-PMC_LODET
C1:PSL-PMC_PZT
C1:PSL-PMC_MODET
3123 - ADC (failed)
C1:PSL-126MOPA_AMPMON
C1:PSL-126MOPA_126MON
C1:PSL-FSS_RCTRANSPD
C1:PSL-FSS_MINCOMEAS
C1:PSL-FSS_RMTEMP
C1:PSL-FSS_RCTEMP
C1:PSL-FSS_MIXERM
C1:PSL-FSS_SLOWM
C1:PSL-FSS_TIDALINPUT
C1:PSL-PMC_PMCTRANSPD
C1:PSL-PMC_PMCERR
C1:PSL-PPKTP_TEMP
4116 - DAC
C1:PSL-126MOPA_126CURADJ
C1:PSL-126MOPA_DCAMP
C1:PSL-126MOPA_DCAMP-
C1:PSL-FSS_INOFFSET
C1:PSL-FSS_MGAIN
C1:PSL-FSS_FASTGAIN
C1:PSL-FSS_PHCON
C1:PSL-FSS_RFADJ
C1:PSL-FSS_SLOWDC
C1:PSL-FSS_VCOMODLEVEL
C1:PSL-FSS_TIDAL
C1:PSL-FSS_TIDALSET
C1:PSL-PMC_GAIN
C1:PSL-PMC_INOFFSET
C1:PSL-PMC_PHCON
C1:PSL-PMC_RFADJ
C1:PSL-PMC_RAMP
XVME-210 - Binary Input
C1:PSL-126MOPA_FAULT
C1:PSL-126MOPA_INTERLOCK
C1:PSL-126MOPA_SHUTTER
C1:PSL-126MOPA_126LASE
C1:PSL-126MOPA_AMPON
XVME-220 - Binary Output
C1:PSL-126MOPA_126NE
C1:PSL-126MOPA_126STANDBY
C1:PSL-126MOPA_SHUTOPENEX
C1:PSL-126MOPA_STANDBY
C1:PSL-FSS_SW1
C1:PSL-FSS_SW2
C1:PSL-FSS_FASTSWEEP
C1:PSL-FSS_PHFLIP
C1:PSL-FSS_VCOTESTSW
C1:PSL-FSS_VCOWIDESW
C1:PSL-PMC_SW1
C1:PSL-PMC_SW2
C1:PSL-PMC_PHFLIP
C1:PSL-PMC_BLANK |
13741
|
Mon Apr 9 18:46:03 2018 |
gautam | Update | IOO | Further debugging |
- I analyzed the data from the free swinging MC test conducted over the weekend. Attachment #1 shows the spectra. Color scheme is same for all panels.
- I am suspicious of MC3: why does the LR coil see almost no Yaw motion?
- The "equilibrium" values of all the sensor signals (at the IN1 of the coil input filters) are within 20% of each other (for MC3, but also MC1 and MC2).
- The position resonance is also sensed more by the side coil than by the LR coil.
- To rule out satellite box shenanigans, I just switched the SRM and MC3 satellite boxes. But coherence between frequency noise as sensed by the arms remain.
- I decided to clean up my IMC nosie budget a bit more.
- Attachment #2 shows the NB as of today. I'll choose a better color palette for the next update.
- "Seismic" trace is estimated using the 40m gwinc file - the MC2 stack is probably different from the others and so it's contribution is probably more, but I think this will suffice for a first estimate.
- "RAM" trace is measured at the CM board input, with MC2 misaligned.
- The unaccounted noise is evident from above ~8 Hz.
- More noises will be added as they are measured.
- I am going to spend some time working on modeling the CM board noise and TF in LTspice. I tried getting a measurement of the transfer function fron IN1 to the FAST output of the CM board with the SR785 (motivation being to add the contribution of the input referred CM board noise to the NB plot), but I suspect I screwed up something w.r.t. the excitation amplitude, as I am getting a totally nonsensical shape, which also seems to depend on my input excitation amplitude. I don't think the output is saturated (viewed during measurement on a scope), but perhaps there are some subtle effects going on.
|
Attachment 1: MC_Freeswinging.pdf
|
|
Attachment 2: IMC_NB_20180409.pdf
|
|
13740
|
Mon Apr 9 16:30:21 2018 |
Kira | Update | PEM | MEDM setup |
I created an MEDM screen for the PID control. In addition, I added a new EPICS channel for the setpoint so that it could be adjusted using the MEDM screen.
Edit: forgot to mention the channel name is C1:PEM-SEIS_EX_TEMP_SETPOINT
Edit #2: the path for the MEDM is /opt/rtcds/caltech/c1/medm/c1pem/C1PEM_SEIS_EX_TCTRL.adl |
Attachment 1: MEDM_screen.png
|
|
13739
|
Mon Apr 9 08:39:39 2018 |
Steve | Update | PEM | M5.3 eq Souther CA |
Earth quake M5.3 2018-04-05 19:29:16UTC Santa Cruz Island, CA
|
Attachment 1: M5.3_Santa_Cruz_Is.CA.png
|
|
Attachment 2: after_M5.3.png
|
|
Attachment 3: M5.3vac.png
|
|
13738
|
Fri Apr 6 22:23:53 2018 |
Kevin | Update | IOO | Coil driver noise |
Quote: |
Why is MC2 LR so different from the others???
|
The previous measurements were made from the coil driver output monitors. To investigate why the MC2 LR coil has less noise than the other coils, I also measured the noise at the output to the coils.
The circuit diagram for the coil driver board is given in D010001 and a picture of the rack is on the 40m wiki here. The coil driver boards are in the upper left quadrant of the rack. The input to the board is the column of LEMOs which are the "Coil Test In" inputs on the schematic. The output monitors are the row of LEMOs to the right of the input LEMOs and are the "FP Coil Volt Mon" outputs on the schematic. The output to the coils "Coil Out" in the schematic are carried through a DB15 connector.
The attachment shows the voltage noise for the MC2 LR coil as well as the UL coil which is similar to all of the other coils measured in the previous measurement. While the LR coil is less noisy than the UL coil as measured at the output monitor, they have the same noise spectrum as measured at the output to the coils themselves. So there must be something wrong with the buffer circuit for the MC2 LR voltage monitor, but the output to the coils themselves is the same as for the other coils. |
Attachment 1: MC2_coil_driver.pdf
|
|
13737
|
Fri Apr 6 21:39:09 2018 |
gautam | Update | IOO | MC2 suspension health checkup |
While Kevin is working on the MC2 electronics chain - we disconnected the output to the optic (DB15 connector on coil driver board). I decided to look at the 'free' freeswinging MC2 OSEM shadow sensor data. Attachment #1 suggests that the suspension eigenmodes are showing up in the shadow sensors, but the 0.8Hz peak seems rather small, especially compared to the result presented in this elog.
Maybe I'll kick all 3 MC optics tonight and let them ringdown overnight, may not be a bad idea to checkup on the health of the MC suspensions/satellite boxes... [MC suspensions were kicked @1207113574]. PSL shutter will remain closed overnight...
Quote: |
Why is MC2 LR so different from the others???
|
|
Attachment 1: MC2_Freeswinging.pdf
|
|
13736
|
Fri Apr 6 18:28:57 2018 |
rana | Update | PEM | PID tuning |
Made some changes:
- Set P and D gains to zero. We only need slow drift control.
- Changed names of the python script and .ini file to distinguish it from the FSS stuff. Lives in scripts/PEM/
- removed debug flag from argParse. To run in non-debug mode you use the "-O" option of python as usual.
- Fixed the upper/lower limit convention for the heater. Was backwards.
- Removed the "rail" function that was defined. We can just use numpy.clip since that's already built in.
There is also now a StripTool file in the scripts/PEM directory which has appropriate channel names and scales for PID loop tuning. Use this file!
I'm leaving it running over the weekend with K_I = -0.003. There is a StripTool on rossa which you can watch. The code itself is running on a tmux session on megatron. Let's ONLY run this code there until we're satisfied that things are good.
Update Sun Apr 8 00:40:11 2018: Lowered gain by factors of 3 down to -0.0001 Saturday afternoon. Seems like still oscillating a bit, now with a ~4 hour period. Setting it to -3e-5 now. Usually we have a linear feedback loop, but our actuation voltage actually gets squared (P = I^2 R) before being integrated to produce temperature. Wonder if we should think of linearizing the feedback control signal to make the loop act nicer.
Update Sun Apr 8 21:09:48 2018: Set K_I = -1e-5 earlier today. Seems to have stabilized nearly, but temperature swings are still +/- 1 K. Will need to add some proportional feedback (K_P) to increase the loop bandwidth, but system is at least sort of stable now. Probably should start construction of EY,BS systems now. |
Attachment 1: HeaterTest.png
|
|
13735
|
Fri Apr 6 16:17:20 2018 |
Kira | Update | PEM | PID tuning |
I have been trying to tune the PID and have managed to descrease the oscillations without saturating the actuator. I'm going to model the system to calculate the exact values of P, I and D in order to get rid of the oscillations altogether. I was going to record the data using Data Viewer, but there seems to be some issue with that, so I'm using StripTool for now. |
Attachment 1: PID_tuning_progress.png
|
|
13734
|
Fri Apr 6 16:07:18 2018 |
gautam | Update | CDS | Frequent EPICS dropouts |
Kira informed me that she was having trouble accessing past data for her PID tuning tests. Looking at the last day of data, it looks like there are frequent EPICS data dropouts, each up to a few hours. Observations (see Attachment #1 for evidence):
- Problem seems to be with writing these EPICS channels to frames, as the StripTool traces do not show the same discontinuities.
- Problem does not seem local to c1auxex (new Acromag machine). These dropouts are also happening in other EPICS channel records. I have plotted PMC REFL, which is logged by the slow machine c1psl, and you can see the dropouts happen at the same times.
- Problem does not seem to happen to the fast channel DQ records (see ETMX Sensor record plotted for 24 hours, there are no discontinuties).
It is difficult to diagnose how long this has been going on for, as once you start pulling longer stretches of data on dataviewer, any "data freezes" are washed out in the extent of data plotted. |
Attachment 1: EPICS_dropout.png
|
|
13733
|
Fri Apr 6 10:00:29 2018 |
gautam | Update | CDS | CDS puzzle |
Quote: |
Clearly, there is a discrepancy for f>20Hz. Why?
|
Spectral leakage |
13732
|
Thu Apr 5 19:31:17 2018 |
gautam | Update | CDS | EPICS processes for c1ioo dead |
I found all the EPICS channels for the model c1ioo on the FE c1ioo to be blank just now. The realtime model itself seemed to be running fine, judging by the IMC alignment (as the WFS loops seemed to still be running okay). I couldn't find any useful info in demsg but I don't know what I'm looking for. So my guess is that somehow the EPICS process for that model died. Unclear why. |
13731
|
Thu Apr 5 13:46:42 2018 |
gautam | Update | SUS | big earthquake |
Seems like there was a 5.3 magnitude EQ ~10km from us (though I didn't feel it). All watchdogs were tripped so our mirrors definitely felt it. ITMX is stuck (but all the other optics are damping fine). I tried the usual jiggling of DC bias voltage but ITMX still seems stuck. Probably a good sign that the magnet hasn't come off, but not ideal that I can't shake it free..
edit: after a bit more vigorous shaking, ITMX was freed. I had to move the bias slider by +/-10,000 cts, whereas initially I was trying +/-2000 cts. There is a tendency for the optic to get stuck again once it has been freed (while the optic's free swinging motion damps out), so I had to keep an eye out and as soon as the optic was freed, I re-engaged the damping servos to damp out the optic motion quickly. |
Attachment 1: EQ_April52018.png
|
|
13730
|
Thu Apr 5 12:13:18 2018 |
Koji | Update | IOO | Coil driver noise |
Why is MC2 LR so different from the others??? |
13729
|
Thu Apr 5 10:38:38 2018 |
gautam | Update | CDS | CDS puzzle |
I'm probably doing something stupid - but I've not been able to figure this out. In the MC1 and MC3 coil driver filter banks, we have a digital "28HzELP" filter module in FM9. Attachment #1 shows the MC1 filterbanks. In the shown configuration, I would expect the only difference between the "IN1" and "OUT" testpoints to be the transfer function of said ELP filter, after all, it is just a bunch of multiplications by filter coefficients. But yesterday looking at some DTT traces, their shapes looked suspicious. So today, I did the analysis entirely offline (motivation being to rule out DTT weirdness) using scipy's welch. Attachment #2 shows the ASDs of the IN1 and OUT testpoint data (collected for 30s, fft length is set to 2 seconds, and hanning window from scipy is used). I've also plotted the "expected" spectral shape, by loading the sos coefficients from the filter file and using scipy to compute the transfer function.
Clearly, there is a discrepancy for f>20Hz. Why?
Code used to generate this plot (and also a datafile to facilitate offline plotting) is attached in the tarball Attachment #3. Note that I am using a function from my Noise Budget repo to read in the Foton filter file...
*ChrisW suggested ruling out spectral leakage. I re-ran the script with (i) 180 seconds of data (ii) fft length of 15 seconds and (iii) blackman-harris window instead of Hanning. Attachment #4 shows similar discrepancy between expectation and measurement... |
Attachment 1: MC1_outputs.png
|
|
Attachment 2: EllipTFCheck_MC1.pdf
|
|
Attachment 3: MC1_ELP.tgz
|
Attachment 4: EllipTFCheck_MC1.pdf
|
|
13728
|
Thu Apr 5 04:36:56 2018 |
Kevin | Update | IOO | Coil driver noise |
[Gautam, Kevin]
We measured the MC coil driver noise at the output monitors of the coil driver board with an SR785 in order to further diagnose the excess IMC frequency noise.
Attachments 1 and 2 show the noise for the UL coils of MC3 and MC2 with various combinations of output filters engaged. When the 28 Hz elliptic filter is on, the analog dewhitening filter is off, and vice versa. The effect of the analog low pass filter is visible in MC3, but the effect of the digital low pass filter is swamped by the DAC noise.
We locked the arms and measured the ALS beatnote in each of these filter combinations, but which filters were on did not effect the excess IMC frequency noise. This suggests that the coil drivers are not responsible for the excess noise.
Attachment 2 shows the noise for all five coils on MC1, MC2, and MC3 as well as for ITMY, which is on a different DAC card from the MCs. All filters were on for these measurements.
|
Attachment 1: MC3.pdf
|
|
Attachment 2: MC2.pdf
|
|
Attachment 3: CoilDriver.pdf
|
|
13727
|
Wed Apr 4 16:23:39 2018 |
gautam | Update | CDS | slow machine bootfest |
[johannes, gautam]
It's been a while - but today, all slow machines (with the exception of c1auxex) were un-telnetable. c1psl, c1iool0, c1susaux, c1iscaux1, c1iscaux2, c1aux and c1auxey were rebooted. Usual satellite box unplugging was done to avoid ITMX getting stuck. |
13726
|
Wed Apr 4 16:23:10 2018 |
Kira | Update | PEM | PID test |
I did a step response for the loop from 35 degrees to 40 degrees. The PID is not properly tuned, so the signal oscillates. In the graph, the blue curve is the temperature of the can in celcius and the green curve is the heating power in watts. The x-axis is in minutes. Before, the signal was too noisy to do a proper step response, so I placed a 3.3 microF capacitor in parallel with the resistor in my temperature sensor circuit (I'll draw and attach this updated version). This created a 5 Hz low pass filter and the signal is now pretty clean.
-----
I also added in new Epics channels so that we could log the data using Data Viewer. The channels I added were C1:PEM-SEIS_EX_TEMP_MON_CELCIUS and C1:PEM-SEIS_EX_TEMP_CTRL_WATTS. I used 13023 as a guide on how to do this.
Update: the channels work and show data in Data Viewer
-----
Edit: I've attached a photo of the circuit with the capacitor indicated. It is in parallel with the resistor below it. I've attached an updated circuit diagram as well. |
Attachment 1: step_response.png
|
|
Attachment 2: capacitor.jpg
|
|
Attachment 3: IMG_20180412_120427.jpg
|
|
13725
|
Mon Apr 2 15:14:21 2018 |
Koji | Update | General | Modulation depth measurement for an aLIGO EOM |
The new matching circuit was tested.
Results:
f_nominal f_actual response required mod. drivng power
[MHz] [MHz] [mrad/V] [rad] needed [dBm]
9.1 9.1 55 0.22 => 22
118.3 118.2 16 0.01 => 6
45.5 45.4 45 0.28 => 25
24.1 N/A 2.1 0.014 => 27
Comments:
- 9.1MHz and 118.3MHz: They are just fine.
- 24.1MHz: Definitely better (>x3) than the previous trial to combine 118MHz & 24MHz.
We got about the same modulation with the 50Ohm terminated bare crystal (for the port1).
So, this is sort of the best we can do for the 24.1MHz with the current approach.
The driving power of 27dBm is required at 24.1MHz
- About the 45MHz
- The driving power of 27dBm is required at 24.1MHz
- The maximum driving power with the AM stabilized driver is 23dBm, nominally to say.
- I wonder how we can reduce resistance (and capacitance) of the 45MHz further...?
- I also wonder if the IFO can be locked with reduced modulation (0.28 rad->0.2 rad)
- Can the driver max power be boosted a bit? (i.e. adding an attenuator in the RF power detection path)
|
Attachment 1: modulation_depth.pdf
|
|
Attachment 2: impedance_eom.pdf
|
|
13724
|
Fri Mar 30 22:37:36 2018 |
Kevin | Update | IOO | MCREFL_PD Optical response measurement |
[Gautam, Kevin]
We redid the measurement measuring the voltage noise from the REFL PD demod board output monitor with an SR785 with the noise eater on and off. We used a 100x preamp to amplify the signal above the SR785 noise. The SR785 noise floor was measured with the input to the preamp terminated with 50 ohms. The spectra shown have been corrected for the 100x amplification.
This measurement shows no difference with the noise eater on or off.
Quote: |
the noise eater on/off measurements should be done for 0-100 kHz and from the demod board output monitor
|
|
Attachment 1: REFLPD_DemodBoard.pdf
|
|
13723
|
Fri Mar 30 16:10:46 2018 |
Kira | Update | PEM | PID test |
I created two new channels today, C1:PEM-SEIS_EX_TEMP_MON_CELCIUS, which turns the output voltage signal into degrees C, and C1:PEM-SEIS_EX_TEMP_CTRL_WATTS, which takes the input voltage from the DAC and turns it into a value of watts. I'm trying to stabilize the temperature at 35 degrees, but it's taking a lot longer than expected. Perhaps we'll need to use different values for P and I and decrease the noise in the sensor, since right now there's about a 10 degree variation between the highest and lowest values. |
13722
|
Fri Mar 30 06:16:45 2018 |
rana | Update | PEM | PID test |
Can't really figure out what this plot means. We need to see the sensor (in units of deg C) and the control signal (in heating power (W)). The plot should show a few step responses with the PID loop on, so that we can see the loop response time. Please zoom in on the axes so that we can see what's happening.
Quote: |
[Kira, Gautam]
We closed the loop today and implemented the PID script. I have attached the StripTool graph for an integral value of 0.5 and proportional value of 20. We had some issues getting it to work properly and it would oscillate between some low values of the control voltage. The set point here was -3.20, which corresponds to about a 20 degree increase in temperature. The next step would be to find which values of Kp, Ki, and Kd would work in this case and low pass filter the signal from the temperature sensor, and also create an MEDM screen for easier PID control.
|
|
13721
|
Fri Mar 30 06:14:31 2018 |
rana | Update | IOO | MCREFL_PD Optical response measurement |
the noise eater on/off measurements should be done for 0-100 kHz and from the demod board output monitor |
13720
|
Fri Mar 30 03:23:50 2018 |
Koji | Update | General | aLIGO EOM work |
I have been working on the aux beat setup on the PSL table between 9PM-3AM.
This work involved:
- Turning off the main marconi
- Turning off the freq generation unit (incl IMC modulation)
- Closing the PSL shutter
After the work, these were reverted and the IMC and both arms have been locked. |
13719
|
Thu Mar 29 17:57:36 2018 |
arijit | Update | IOO | MCREFL_PD Optical response measurement |
Kevin, Gautam and Arijit
Today we performed the in-loop noise measurements of the MCREFL-PD using the SR785 to ascertain the effect of the Noise Eater on the laser. We took the measurements at the demodulated output channel from the MCREFL-PD. We performed a series of 6 measurements with the Noise Eater ''ON'' and ''OFF''. The first data set is an outlier probably, due to some transient effects. The remaining data sets were recorded in succession with a time interval of 5 minutes each between the Noise Eater in the ''ON'' and ''OFF'' state. We used the calibration factor of 13kHz/Vrms from elog 13696 to convert the V_rms to Hz-scale.
The conclusion is that the NOISE EATER does not have any noticeable effect on the noise measurements.
ALS beat spectrum and also the arm control signal look as they did before. coherence between arm control signals (in POX/POY lock) is high between 10-100Hz, so looks like there is still excess frequency noise in the MC transmitted light. Looking at POX as an OOL sensor with the arm under ALS control shows ~10x the noise at 100 Hz compared to the "nominal" level, consistent with what Koji and I observed ~3weeks ago.
We tried swapping out Marconis. Problem persists. So Marconi is not to blame. I wanted to rule this out as in Jan, Steve and I had installed a 10MHz reference to the rear of the Marconi. |
Attachment 1: NOISE_EATER_On_OFF.pdf
|
|
13718
|
Thu Mar 29 17:14:42 2018 |
Kira | Update | PEM | PID test |
[Kira, Gautam]
We closed the loop today and implemented the PID script. I have attached the StripTool graph for an integral value of 0.5 and proportional value of 20. We had some issues getting it to work properly and it would oscillate between some low values of the control voltage. The set point here was -3.20, which corresponds to about a 20 degree increase in temperature. The next step would be to find which values of Kp, Ki, and Kd would work in this case and low pass filter the signal from the temperature sensor, and also create an MEDM screen for easier PID control. |
Attachment 1: PID_test.png
|
|
13717
|
Thu Mar 29 12:03:37 2018 |
Jon Richardson | Summary | General | Proof-of-Concept SRC Gouy Phase Measurement |
I've been developing an idea for making a direct measurement of the SRC Gouy phase at RF. It's a very different approach from what has been tried before. Prior to attempting this at the sites, I'm interested in making a proof-of-concept measurement demonstrating the technique on the 40m. The finesse of the 40m SRC will be slightly higher than at the sites due to its lower-transmission SRM. Thus if this technique does not work at the 40m, it almost certainly will not work at the sites.
The idea is, with the IFO locked in a signal-recycled Michelson configuration (PRM and both ETMs misaligned), to inject an auxiliary laser from the AS port and measure its reflection from the SRC using one of the pre-OMC pickoff RFPDs. At the sites, this auxiliary beam is provided by the newly-installed squeezer laser. Prior to injection, an AM sideband is imprinted on the auxiliary beam using an AOM and polarizer. The sinusoidal AOM drive signal is provided by a network analyzer, which sweeps in frequency across the MHz band and demodulates the PD signal in-phase to make an RF transfer function measurement. At the FSR, there will be a AM transmission resonance (reflection minimum). If HOMs are also present (created by either partially occluding or misaligning the injection beam), they too will generate transmission resonances, but at a frequency shift proportional to the Gouy phase. For the theoretical 19 deg one-way Gouy phase at the sites, this mode spacing is approximately 300 kHz. If the transmission resonances of two or more modes can be simultaneously measured, their frequency separation will provide a direct measurement of the SRC Gouy phase.

The above figure illustrates this measurement configuration. An attached PDF gives more detail and the expected response based on Finesse modeling of this IFO configuration. |
Attachment 1: src_gouy_phase_v3.pdf
|
|
13716
|
Wed Mar 28 21:47:37 2018 |
arijit | Update | IOO | MCREFL_PD Optical response measurement |
Kevin, Gautam and Arijit
We did a optical measurement of the MCREFL_PD transimpedance using the Jenny Laser set-up. We used 0.56mW @1064nm on the NewFocus 1611 Photodiode as reference and 0.475mW @1064nm on the MCREFL_PD. Transfer function was measured using the AG4395 network analyzer. We also fit the data using the refined LISO model. From the optical measurement, we can see that we do not have a prominent peak at about 300MHz like the one we had from the electrical transimpedence measurement. We also put in the electrical transimpedence measurement as reference. RMS contribution of 300MHz peak to follow.
As per Rana`s advice I have updated the entry with information on the LISO fit quality and parameters used. I have put all the relevant files concerning the above measurement as well as the LISO fit and output files as a zip file "LISO_fit" . I also added a note describing what each file represents. I have also updated the plot with fit parameters and errors as in elog 10406. |
Attachment 1: LISO_fit_with_info.pdf
|
|
Attachment 2: LISO_fit.zip
|
13715
|
Wed Mar 28 21:31:39 2018 |
gautam | Update | IOO | MC REFL PD removed |
I re-installed the MC REFL photodiode. Centered beam on the PD by adjusting steering mirror to maximize the DC signal level (on o'scope) at the DC monitoring port. Curiously, the DC level on the scope (high-Z) was ~2.66V DC, whereas the MEDM screen reports ~twice that value, at ~5.44 "V". We may want to fix this "calibration" (or better yet, use physical units like mW). Noise-eater On/Off comparison of MC error signals to follow. |
13714
|
Wed Mar 28 17:28:58 2018 |
Steve | Update | PSL | noise eater on or off |
Till RIN measurement noise eater is off on 2W laser. The inno 1W has no noise eater.
2010 power v current table is below.
Quote: |
Koji and Kevin measured the output power vs injection current for the Innolight 2W laser.
The threshold current is 0.75 A.
The following data was taken with the laser crystal temperature at 25.04ºC (dial setting: 0.12).
Injection Current (A) |
Dial Setting |
Output Power (mW) |
0.000 |
0.0 |
1.2 |
0.744 |
3.66 |
1.1 |
0.753 |
3.72 |
4.6 |
0.851 |
4.22 |
102 |
0.954 |
4.74 |
219 |
1.051 |
5.22 |
355 |
1.151 |
5.71 |
512 |
1.249 |
6.18 |
692 |
1.350 |
6.64 |
901 |
1.451 |
7.08 |
1118 |
1.556 |
7.52 |
1352 |
1.654 |
7.92 |
1546 |
1.761 |
8.32 |
1720 |
1.853 |
8.67 |
1855 |
1.959 |
9.05 |
1989 |
2.098 |
9.50 |
2146 |
|
|
Attachment 1: inno2W.png
|
|
Attachment 2: inno1W.png
|
|
13713
|
Wed Mar 28 16:44:27 2018 |
Steve | Update | General | AP table today |
MCRefl is absent, it is under investigation. I removed a bunch of hardware and note all spare optics along the edges.
|
Attachment 1: AP_Table_20180328.png
|
|
13712
|
Tue Mar 27 23:37:35 2018 |
gautam | Update | IOO | MC REFL PD removed |
I've removed the MC REFL PD unit from the AS table for investigation. So MC won't lock.
PSL shutter was closed and location of PD was marked with sharpie (placing guides to indicate position wasn't convenient). I also kapton taped the PD to minimize dust settling on the PD while I have it out in the electronics area. Johannes has the camera, and my cellphone image probably isn't really high-res enough for diagnostics but I'm posting it here anyways for what it's worth. More importantly - the board is a D980454 revision B judging by the board, but there is no schematic for this revision on the DCC. The closest I can find is a D980454 Rev D. But I can already see several differences in the component layout (though not all of them may be important). Making a marked up schematic is going to be a pain . I'm also not sure what the specific make of the PD installed is.
The lid of the RF cage wasn't on.
More to follow tomorrow, PD is on the electronics workmench for now...
gautam 28 March 2018: Schematic has been found from secret Dale stash (which exists in addition to the secret Jay stash). It has also been added to the 40m electronics tree. |
Attachment 1: IMG_6955.JPG
|
|
13711
|
Tue Mar 27 19:32:03 2018 |
arijit | Update | IOO | PSL noise eater was off |
Kevin, Gautam and Arijit
We made a measurement of the MC_REFL photodiode transfer function using the network analyzer. We did it for two different power input (0dB and -10dB) to the test measurement point of the MC_REFL photodiode. This was important to ensure the measurements of the transfer function of the MC_REFL photodiode was in the linear regime. The measurements are shown in attachment 1. We corrected for phase noise for the length of cable (50cm) used for the measurement. With reference to ELOG 10406, in comparison to the transimpedance measurement performed by Riju and Koji, there is a much stronger peak around 290MHz as observed by our measurement.
We also did a noise measurement for the MC_REFL photodiode. We did it for three scenarios: 1. Without any light falling on the photodiode 2. With light falling on the photodiode, the MC misaligned and the NPRO noise eater was OFF 3. With light falling on the photodiode, the MC misaligned and the NPRO noise eater was ON. We observed that the noise eater does reduce the noise being observed from 80kHz to 20MHz. This is shown in attachment 2.
We did the noise modelling of the MC_REFL photodiode using LISO and tried matching the expected noise from the model with the noise measurements we made earlier. The modeled noise is in good agreement with the measured noise with 10Ohms in the output resistance. The schematic for the MC_REFL photodiode however reveals a 50Ohm resistance being used. The measured noise shows excess noise ~ 290MHz. This is not predicted from the simplied LISO model of the photodiode we took.
Discussion with Koji and Gautam revealed that we do not have the exact circuit diagram for the MC_REFL photodiode. Hence the simplified model that was assumed earlier is not able to predict the excess noise at high frequencies. One thing to note however, is that the excess noise is measured with the same amplitude even with no light falling on the MC_REFL photodiode. This means that there is a positive feedback and oscillation in the op-amp (MAX4107) at high frequencies. One way to refine the LISO model would be to physically examine the photodiode circuit.
We also recorded the POX and POY RF monitor photodiode outputs when the interferometer arms are independently stabilized to the laser. Given the noise outputs from the RF photodiodes were similar, we have only plotted the POY RF monitor output for the sake of clarity and convenience.
Quote: |
While Kevin and Arijit were doing their MC_REFL PD noise measurements (which they will elog about separately shortly), I noticed a feature around 600kHz that reminded me of the NPRO noise eater feature. This is supposed to suppress the relaxation oscillation induced peak in the RIN of the PSL. Surprisingly, the noise eater switch on the NPRO front panel was set to "OFF". Is this the normal operating state? I thought we want the noise eater to be "ON"? Have to measure the RIN on the PSL table itself with one of the many available pick off PDs. In any case, as Attachment #1 showed, turning the noise eater back on did not improve the excess IMC frequency noise.
|
|
Attachment 1: MCREFL_TF.pdf
|
|
Attachment 2: MCREFL_SPECTRUM.pdf
|
|
13710
|
Tue Mar 27 11:11:16 2018 |
Kira | Update | PEM | Channel setup |
[Kira, Gautam]
We setup the channels for PID control of the seismometer can. First, we ssh into c1auxex and went to /cvs/cds/caltech/target/c1auxex2 and found ETMXaux.db. We then added in new soft channels that we named C1:PEM-SEIS_EX_TEMP_SLOWKP, C1:PEM-SEIS_EX_TEMP_SLOWKI, C1:PEM-SEIS_EX_TEMP_SLOWKD that will control the proportional, integral and differential gain respectively. These channels are used in the script FSSSlow.py for PID control. We then had to restart the system, but first we turned off the LSC mode and then shut down the watchdog on the X end. After doing the restart, we disabled the OPLEV as well before restarting the watchdog. Then, we enabled the LSC mode again. This is done to not damage any of the optics during the restart. The restart is done by using sudo systemctl restart modbusIOC.service and restarted with sudo systemctl status modbusIOC.service. Then, we made sure that the channels existed and could be read and writtten to, so we tried z read [channel name] and it read 0.0. We then did z write [channel name] 5, and it wrote 5 to that channel. Now that the channels work, we can implement the PID script and check to make sure that it works as well. |
13709
|
Tue Mar 27 08:58:21 2018 |
Steve | Update | VAC | VM1 opened.......scan fine |
Quote: |
CC1 old MKS cold cathode gauge randomly turns on- off. This makes software interlock close VM1 to protect RGA So the closed off RGA region pressure goes up and the result is distorted RGA scan.
CC1 MKS gauge is disconnected and VM1 opened. This reminds me that we should connect our interlocks to CC1 Hornet Pressure gauge.
Quote: |
Pumpdown 80 at 511 days and pd80b at 218 days
Valve configuration: special vacuum normal, annuloses are not pumped at 3 Torr, IFO pressure 7.4e-6 Torr at vac envelope temp 22 +- 1C degrres
|
|
|
Attachment 1: rga2018march27.png
|
|
13708
|
Tue Mar 27 01:39:44 2018 |
gautam | Update | IOO | PSL noise eater was off |
While Kevin and Arijit were doing their MC_REFL PD noise measurements (which they will elog about separately shortly), I noticed a feature around 600kHz that reminded me of the NPRO noise eater feature. This is supposed to suppress the relaxation oscillation induced peak in the RIN of the PSL. Surprisingly, the noise eater switch on the NPRO front panel was set to "OFF". Is this the normal operating state? I thought we want the noise eater to be "ON"? Have to measure the RIN on the PSL table itself with one of the many available pick off PDs. In any case, as Attachment #1 showed, turning the noise eater back on did not improve the excess IMC frequency noise. |
Attachment 1: IR_ALS_20180326.pdf
|
|
13707
|
Mon Mar 26 23:49:27 2018 |
gautam | Update | General | New ADC Adaptor Board installed in C1LSC expansion chassis |
Todd informed me that the ADC Timing adaptor boards we had ordered arrived today. I had to solder on the components and connectors as per the schematic, though the main labor was in soldering the high density connectors. I then proceeded to shut down all models on c1lsc (and then the FE itself). Then classic problem of all vertex machines crashing when unloading models on c1lsc happened (actually Koji noticed that this was happening even on c1ioo). Anyways this was nothing new so I decided to push ahead.
I had to get a cable from Downs that connects the actual GS ADC card to this adaptor board. I powered off the expansion chassis, installed the adaptor board, connected it to the ADC card and restarted the expansion chassis and also the FE. I also reconnected the SCSI cable from the AA board to the adaptor card. It was a bit of a struggle to get all the models back up and running again, but everything eventually came back(after a few rounds of hard rebooting). I then edited the c1x04 and c1lsc simulink models to reflect the new path for the X arm ALS error signals. Seems to work alright.
At some point in the afternoon, I noticed a burning smell concentrated near the PSL table. Koji traced the smell down to the c1lsc expansion chassis. We immediately powered the chassis off. But Steve later informed me that he had already noticed an odd burning smell in the morning, before I had done any work at the LSC rack. Looking at the newly installed adaptor card, there wasn't any visual evidence of burning. So I decided to push ahead and try and reboot all models. Everything came back up normally eventually, see Attachment #1. Particle count in the lab seems a little higher than usual (actually, according to my midnight measurement, they are ~factor of 10 lower than Steve's 8am measurements), but Steve didn't seem to think we should read too much into this. Let's monitor the situation over the coming days, Steve should comment on the large variance seen in the particle counter output which seems to span 2 orders of magnitude depending on the time of the day the measurement is made... Also note that there is a BIO card in the C1LSC expansion chassis that is powered by a lab power supply unit. It draws 0 current, even though the label on it says otherwise. I a not sure if the observed current draw is in line with expectations.
The spare (unstuffed) adaptor cards we ordered, along with the necessary hardware to stuff them, are in the Digital FE hardware cabinet along the east arm.
Steve: particle count in the 40m is following outside count, wind direction, weather condition .....etc. The lab particle count is NOT logged ! This is bad practice. |
Attachment 1: CDS_20180326.png
|
|
13706
|
Mon Mar 26 21:40:26 2018 |
gautam | Summary | IOO | MC2 classical radiation pressure noise |
[rana, gautam]
we measured the RIN of the MC2 transmission using the PDA255 I had put on the MC2 trans table sometime ago for ringdowns. Attached are (i) spectra for the RIN, (ii) spectra for the classical rad. pressure noise assuming 500W circulating power and (iii) a tarball of data and code used to generate these plots.
We took a full span measurement (to make sure there aren't any funky high-freq features) and a measurement from DC-800 Hz (where we are looking for excess noise). The DC level of light on the photodiode was 2.76V (measured using o'scope)
I'll add this to the noise budget later. But the measured RIN seems consistent with a 2013 measurement at 100Hz (though the 2013 measurement is using DTT and so doesn't have high frequency information). |
Attachment 1: IMC_RIN.pdf
|
|
Attachment 2: IMC_RadPress.pdf
|
|
Attachment 3: MC2_radPress.tgz
|
13705
|
Mon Mar 26 21:25:55 2018 |
rana | Summary | IOO | MC2 Trans table has issues |
Gautam, Rana
While at the MC2 table, we noticed that it has some optical problems:
- There is an ND filter mounted to the beam reducing lens. ND filters are illegal, Steve. It causes too much scattering noise. We should instead have a beamsplitter and a dump.
- One of those bad U100-AC28 mounts is in use. This is one of those ones with plastic clips that Osamu liked, but the plastic gets in the way of the beam. Needs to be removed.
- Reflections from the QPD and the PDA255 are not dumped. This causes noise. Bad.
- The QPD transimpedance should be reduced so that it can handle more light. I don't know what it has now, but its probably 10-100 kOhm.
We estimated that the power in the IMC is (1 W)*Finesse/pi = 500 W. The MC2 Transmission spec is < 10 ppm, so the power on the table is probably ~5 mW. Since the PDA255 has a transimpedance of 10 kOhm and a max output power of 10V, it can handle up to ~1 mW. Probably we can get the QPD to handle 4 mW.
Gautam, Steve 3-27
We measured MC2 transmitted power right at the uncoated window ~2.5 mW The beam was just a little bigger than the meter. |
Attachment 1: 20180326_201929.jpg
|
|
13704
|
Mon Mar 26 16:10:33 2018 |
Kira | Update | PEM | final setup sketch |
I made sketches of the final setup. There will be a box in the rack that contains both the heater circuit and the temperature sensor boards. One of them is in the loop while the other isn't. Instead of having many cables leading to the can, there will only be these three, though they can be made into a single wire. It will be connected to the can through a D-9 connector. The second attachment is what will be inside of the box, with all the major wires and components labeled.
-----
Edit: I've canged the layout to (hopefully) make the labels easier to read. I've also added in a cable to the ADC that reads out the voltage across the 1 ohm resistor. I also attached the circuit diagrams for the heater circuit and the temperature sensors. The one for the heater circuit was made by Kevin and I used the same design, except I have LM7818 and LM7918, since the 15V ones were not available at the time I made the circuit.
In addition, all the wires leading to the can will all be part of one bundle of wires (I didn't clearly indicate it as such). There will be a total of 6 wires: two are needed for the wire to supply power to the heater and will have a LEMO connector on the rack end and two are needed for each temperature sensor, which will be attached to the board directly on the rack end.
Also, we don't need two voltage regulators for each temperature circuit. We can just have one of each of LM7815 and LM7915 to supply +/- 15V to the boards. |
Attachment 1: heater_1_new.png
|
|
Attachment 2: heater_2_new.png
|
|
Attachment 3: HeaterCircuit.pdf
|
|
Attachment 4: temp_sensor.png
|
|
13703
|
Mon Mar 26 10:15:20 2018 |
Arijit | Update | IOO | PMC and IMC re-locked |
[Gautam, Arijit]
PMC and IMC re-aligned and re-locked. Both cavities are staying locked. Arm cavities are also locked. |
13702
|
Mon Mar 26 09:25:18 2018 |
Steve | Update | VAC | VM1 opened |
CC1 old MKS cold cathode gauge randomly turns on- off. This makes software interlock close VM1 to protect RGA So the closed off RGA region pressure goes up and the result is distorted RGA scan.
CC1 MKS gauge is disconnected and VM1 opened. This reminds me that we should connect our interlocks to CC1 Hornet Pressure gauge.
Quote: |
Pumpdown 80 at 511 days and pd80b at 218 days
Valve configuration: special vacuum normal, annuloses are not pumped at 3 Torr, IFO pressure 7.4e-6 Torr at vac envelope temp 22 +- 1C degrres
|
|
Attachment 1: CC4VM1.png
|
|
13701
|
Fri Mar 23 12:45:08 2018 |
Kira | Update | PEM | test setup |
I fit the data that we got from the test. The time constant for the cooling came out to be about 4.5 hours. The error is quite large and we should add a low pass filter to the temperature sensor eventually in order to minimize the noise of the measurements. |
Attachment 1: seis_fit.png
|
|
13700
|
Fri Mar 23 12:00:20 2018 |
rana | Update | PEM | test setup |
we don't ever want to use our 16 kHz real time system for such low frequency action; its main purpose is for real-time controls, whereas we are OK with multiple seconds of delay in a thermal loop. The Python PID script is sufficient and highly reliable (after years of testing). |